OpenStudy (lilday12):

Simplify the variable expression by evaluating its numerical part, and enter your answer below. j + 16 - 148 + 4

alones (alones):

Do you need this simplified or evaluated

alones (alones):

if your simplifiying then j+16−148+4 =j+16+−148+4 Combine Like Terms: =j+16+−148+4 =(j)+(16+−148+4) =j+−128 =j−128

OpenStudy (cmtboy2):

j=(-128)

OpenStudy (cmtboy2):

-128+16+-148+4=-256